I need to know what the excluded factor of 20a+4 over 28a is
Answer by stanbon(75887)

20a+4 over 28a is
----------
(4(5a+1))/(28a) = (4/4)[(5a+1)/(7a)] = (5a+1)/(7a)
--------------------
Cheers,
Stan H.
-------------------